OpenStudy (ggrree):

Integral of tan^2(x)sec(x)

OpenStudy (bahrom7893):

u is Sec(x)!!!!

OpenStudy (bahrom7893):

du is tan^2(x) dx.. the rest is for myin.

OpenStudy (turingtest):

not quite bahrom

myininaya (myininaya):

\[\int\limits_{}^{}\tan(x) \cdot \tan(x) \sec(x) dx=\tan(x) \cdot \sec(x)-\int\limits_{}^{}\sec^2(x) \cdot \sec(x) dx\]

OpenStudy (bahrom7893):

oh wait lol i got it the other way around hahaha

myininaya (myininaya):

And you can look at the integral sec^3(x) and use integration by parts there

myininaya (myininaya):

oh wait i didn't need to do what i did

myininaya (myininaya):

\[\int\limits_{}^{}(\sec^2(x)-1)\sec(x) dx=\int\limits_{}^{}\sec^3(x) dx-\int\limits_{}^{}\sec(x) dx\]

OpenStudy (turingtest):

that person is leaving, just so you know they told me on another post

myininaya (myininaya):

\[\int\limits_{}^{}\sec^3(x) dx=\int\limits_{}^{}\sec^2(x) \sec(x)dx=\tan(x) \sec(x)-\int\limits_{}^{}\tan(x) \sec(x) \tan(x) dx\] \[\tan(x) \sec(x)-\int\limits_{}^{}(\sec^2(x)-1) \sec(x) dx\]

myininaya (myininaya):

:(

OpenStudy (turingtest):

I was doing problems with them for quite a while, but don't worry pretty sure they got what you were saying

myininaya (myininaya):

yay i think there is enough info here

OpenStudy (ggrree):

thanks to all of you! I got it now.